OpenStudy (anonymous):

The electrical resistance R (in ohms) of a wire is directly proportional to its length L (in feet). If 250 feet of wire has a resistance of 1.3 ohms, find the resistance for 180 ft of wire

OpenStudy (ybarrap):

R = k L k = R/L = 1.3/250 So R(180) = k180 = (1.3/250)*180 = 0.936 ohms Ans. 0.936 ohms